Let f(x) = px + q, where p and q are real numbers. Find $p+q$ if $(f(f(x))) = 8x + 21.
Found 2 solutions by MathLover1, MathTherapy:
Answer by MathLover1(20850)
Answer by MathTherapy(10552)

Let f(x) = px + q, where p and q are real numbers. Find $p+q$ if $(f(f(x))) = 8x + 21.
If f(f(f(x))) = 8x + 21, then p + q = 5.
If f(f(x)) = 8x + 21, then that's a different story.